OpenStudy (anonymous):

10m-p=-n, for m show work please

OpenStudy (anonymous):

10m-p=-n p-n/10

OpenStudy (anonymous):

=m

OpenStudy (tcp2315):

10m-p=-n 10m -p (+p) = -n (+p) 10m= -n + p 10m/10 = (-n+p)/10 m= (-n +p)/10 or m= (p-n)/10